The distance traveled by a ball rolling down a ramp is given by s(t)=2.25t^2, where t is the time in seconds after the ball is released, and s(t) is measured in feet. Evaluate the average velocity of the ball for the time interval [1.0,1.1].
Please show me your work in this problem, Thank you!

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

The distance traveled by a ball rolling down a ramp is given by s(t)=2.25t^2, where t is the time in seconds after the ball is released, and s(t) is measured in feet. Evaluate the average velocity of the ball for the time interval [1.0,1.1].
Please show me your work in this problem, Thank you!
-------------
Average Velocity = [s(1.1)-s(1)]/[1.1-1]
===
= [2.25(1.1^2) - 2.25(1^2)]/[0.1]
----
= [2.25(0.21)]/0.1
-----
= 4.725
------------
Cheers,
Stan H.
------------

Answer by rothauserc(4718)

Using given formula calculate the distance covered at t=1.0 and t=1.1, then subtract s(1.0) from s(1.1)
s(1.1) = 2.25*(1.1)^2 = 2.7225
s(1.0) = 2.25*(1.0)^2 = 2.25
s(1.1) - s(1.0) = 0.4725
average velocity(Va) equals (delta distance) / (delta time)
Va = 0.4725 / 0.1 = 4.725